# Predicting the structure of dynamic graphs

### 08 Jan 2024 v1

@@ 3.3 Adapting linear and integer programming to graph prediction
(matrix $\bm{S}$ from graph $\bm{\mathcal{G}}$) For a graph with $n$ vertices, let $S$ denote the $n\times n(n-1)/2$ matrix obtained from the adjacency matrix $A_{n\times n}$ as follows: The columns represent the edges obtained from unrolling the upper triangular part of $A$ row by row, and the rows of $S$ represent the vertices 1 to $n$. If an edge $e_{ij}$ is present, i.e., $e_{ij}=1$ with $i<j$ then $s_{ik}=s_{jk}=1$ where $k$ denotes the column corresponding to edge $e_{ij}$, with $k=\left((n-1)+(n-2)+\ldots+(n-i+1)\right)+(j-i)=\frac{1}{2}(i-1)(2n-i)+(j-i)$.
$S=\begin{bmatrix}1&0&1&0&0&0\\ 1&0&0&1&0&0\\ 0&0&0&1&0&1\\ 0&0&1&0&0&1\\ \end{bmatrix}\,,$
The matrix $S$ for predicting the graph at $T+h$ is obtained by considering the union graph from the time series of graphs and new vertices as per the prediction $\hat{n}_{T+h}$.
Definition 3.2.
($\bm{S}$ from a graph sequence $\bm{\{{\mathcal{G}_{t}}\}_{t=1}^{T}}$) For a sequence of graphs $\{{\mathcal{G}_{t}}\}_{t=1}^{T}$ we consider the graph union $\mathcal{G}_{U}=\mathcal{G}_{1}\bigcup\ldots\bigcup\mathcal{G}_{T}$, and add vertices $\hat{n}_{T+h}-n_{T}$, if this quantity is positive. Furthermore, we add all possible edges from the new vertices to the union graph $\mathcal{G}_{U}$ and compute matrix $S$.
@@
where $\bm{u}$ denotes an $n(n-1)/2$ vector of edges and $\bm{d}$ denotes an $n$ vector of predicted degrees with $n=\hat{n}_{T+h}$. For all new (unseen) vertices $k$ we use the same upper bound $f_{u}\left(\hat{d}_{k,T+h}\right)$ computed by taking a mean over the node degree for new nodes from the graph time series $\{\mathcal{G}_{t}\}_{t=1}^{T}$ for $t\leq T$. That is, for each graph $\mathcal{G}_{t}$ in the time series, we consider the new nodes at time $t$ and take their average degree. For the remainder of this section we use $n$ to denote $\hat{n}_{T+h}$.
Of the many solutions present in this space, we want to select a solution that depicts the most likely graph at time $T+h$. However, the term “most likely” can be subjective. Our definition of most likely involves maximizing the function $\sum_{ij}\xi_{ij}\hat{e}_{ij,T+h}$ where $\xi_{ij}$ denote the coefficients of the optimization function. The coefficients satisfy $0\leq\xi_{ij}\leq 1$ with higher values given to edges more likely to be present in $\mathcal{G}_{T+h}$. We state the complete optimization problem below:
$\max$ $\sum_{i,j\in V_{T+h}}\xi_{ij}\hat{e}_{ij,T+h}\,,$
$\text{with}\qquad\bm{0}$ $\leq S\bm{u}\leq f_{u}\left(\bm{d}\right)\,,$ (3)
@@
    This coefficient function is naive as it considers any edge to be equally likely including edges that have never been present in the past. This function may be useful for instances with limited history or when we want to completely disregard the history of a network when predicting edges at time $T+h$.
@@
    Proportional coefficients (C3): This scheme has a parameter $\lambda$. Here we have
    $\xi_{ij}=\begin{cases}\frac{\sum_{t\leq T}|e_{ij,t}=1|}{\sum_{t\leq T}\left|\left(v_{i}\in\mathcal{G}_{t}\right)\ \cap\left(v_{j}\in\mathcal{G}_{t}\right)\right|}&\text{for existing vertices}\\ \lambda&\text{if }i\text{ or }j\text{ are new vertices}\end{cases}$ (5)
    where $\sum_{t\leq T}|e_{ij,t}=1|$ denotes the number of times edge $e_{ij}$ is present in the graph time series, and $\sum_{t\leq T}\left|\left(v_{i}\in\mathcal{G}_{t}\right)\ \cap\left(v_{j}\in\mathcal{G}_{t}\right)\right|$ denotes the number of times both $v_{i}$ and $v_{j}$ are present. Thus, for existing vertices at time $T$, we take the proportion of time points when edge $e_{ij}$ is present. For new vertices, we assign a parameter $\lambda$, which can be set to a fixed value (such as $0.01$) or a variable taking into account the distribution of $\xi_{ij}$ for existing vertices.
@@
    Predictive coefficients (C4): Here we use a prediction of $e_{ij,T+h}$ as the coefficient. This scheme too involves a parameter $\lambda$.
    $\xi_{ij}=\begin{cases}f\left(e_{ij}\right)&\text{ for existing vertices }i,j\\ \mu_{i}&\text{ if }v_{i}\text{ is existing and }v_{j}\text{ is new }\\ \lambda&\text{if both }v_{i},v_{j}\text{ are new}\end{cases}$ (6)
    where $f\left(e_{ij}\right)$ is a prediction of $e_{ij}$ using the time series $\{e_{ij}\}_{t=1}^{T}$ and $\mu_{i}$ denotes a vertex-based coefficient for existing vertices when connecting with new vertices. For example, a more connected vertex can have a higher coefficient $\mu_{i}$ compared to a less connected vertex.
Of the different coefficient functions, binary and proportional coefficients can be considered as simpler approximations of predictive coefficients. We explore binary and proportional coefficients, but not uniform or predictive coefficients because uniform coefficients are too naive and predictive coefficients are time consuming.
Considering proportional/predictive coefficients $\{\xi_{ij}\}_{v_{i},v_{j}\in V_{t+H}}$, the question arises why these alone cannot be used to describe the graph $\mathcal{G}_{T+h}$. We emphasize that $\xi_{ij}$ is computed by considering individual edges $e_{ij}$ or nodes $v_{i},v_{j}$, without taking the network structure into account. Thus, $\xi_{ij}$ is not a good approximation for $e_{ij,T+h}$ and we do not recommend it to be used by itself to construct $\mathcal{G}_{T+h}$. However, $\xi_{ij}$ provides good value as coefficients of the optimization problem described in equation (3.3). As we later consider modifications to the constraint space, we state the problem formally.
Formulation 1 (F1) We consider the following optimization problem:
@@
$\text{with}\qquad\bm{0}\leq S\bm{u}$ $\leq f_{u}\left(\bm{d}\right)\,,$ (7)
with $\bm{u}=\{u_{\ell}\}_{\ell=1}^{n(n+1)/2}$ and $u_{\ell}\in\{0,1\}$ where each inequality in (3.3) describes a constraint for the degree upper bound of a single vertex $i$ given by
$\sum_{j}\hat{e}_{ij,T+h}\leq f_{u}\left(\hat{d}_{i,T+h}\right)\,.$
We first show the columns of $S$ add to either 2 or 0.
Lemma 3.3.
For unweighted, undirected graphs without loops let $S=\{s_{k\ell}\}_{n\times n(n-1)/2}$. Then for any $k$, $\sum_{k}s_{k\ell}\in\{0,2\}$ , i.e., the column sums of $S$ equals either 0 or 2.
Proof.
The rows of $S$ denote vertices and its columns denote edges. Suppose the $\ell$th column denotes the edge $e_{ij}$. That means in the $\ell$th column
$s_{k\ell}=\begin{cases}1,\quad e_{ij}=1\text{ and }k=i\text{ or }j\\ 0,\quad\text{otherwise}\,\end{cases}\,.$
Thus,
$\sum_{k}s_{k\ell}=\begin{cases}2,\quad e_{ij}=1\\ 0,\quad\text{otherwise}\,\end{cases}\,.$
@@
The proof is given in Appendix A.
So far each constraint considers the degree of a single node. However, as we are considering upper bounds we may add a lot more edges to the graph. By constraining the total edges using its prediction can help us to get a more realistic graph. We note that this is a global constraint affecting all nodes.
Formulation 2 (F2) We consider the following optimization problem:
$\max$ $\sum_{i,j\in V_{T+h}}\xi_{ij}\hat{e}_{ij,T+h}\,,$
$\text{with}\qquad\bm{0}$ $\leq S\bm{u}\leq f_{u}\left(\bm{d}\right)\,,$
3.4 Distribution of graphs
By solving the optimization problem described in Formulations 1 or 2 we obtain a possible graph at time $T+h$, which we denote by $\hat{\mathcal{G}}_{T+h}$. This graph is a prediction and thus is an element of a predictive distribution. In fact it was obtained by making certain choices on $n$, (used to denote $\hat{n}_{T+h}$) and $f_{u}(\bm{d})$.
Figure 1 illustrates the predicted graph distribution in which $\hat{\mathcal{G}}_{T+h}$ is a function of predicted nodes $\hat{n}_{T+h}$ and $f_{u}(\bm{d})$. If we explore the predictions for different values of $\hat{n}_{T+h}$ and different upper bounds $u$, we get different graphs. Let $\gamma$ denote the quantile used to obtain $\hat{n}_{T+h}$ and $u$ denote the quantiles in $f_{u}(\bm{d})$. Then we can think of the graph $\hat{\mathcal{G}}_{T+h}$ as belonging to a graph distribution $\mathfrak{G}$ having parameters $\gamma$ and $u$:
@@ 3.4 Distribution of graphs
While these are explicit parameters, the coefficient scheme used to obtain $\xi_{ij}$, the methods used for predicting individual time series and their hyperparameters can also be considered as part of a wider parameter pool. While we use ARIMA models with automatic parameter selection (Hyndman & Khandakar, 2008) for time series prediction, LSTMs and other prediction methods can also be used.
Figure 1: A simple illustration of the predicted graph distribution in terms of $\gamma$ and $u$. The parameter $\gamma$ gives rise to different $\hat{n}_{T+h}$ values. The 2 graphs on the left have 4 nodes each, corresponding to a single $\gamma$ (or $\hat{n}_{T+h}$) value, but has different number of edges corresponding to different $u$ values. Similarly, the two graphs on the right have 7 nodes corresponding to a higher value of $\gamma$ with the rightmost graph having more edges due to a higher $u$.
3.5 Evaluation Metrics
We use the following evaluation metrics: 1. absolute node error $\frac{|\hat{n}_{T+h}-n_{T+h}|}{n_{T+h}}$ gives the absolute error of the predicted number of vertices compared to the actual as a ratio. 2. The adjacency matrix error computed as $\frac{\lVert\hat{A}_{T+h}-{A}_{T+h}\rVert}{\lVert{A}_{T+h}\rVert}$, where we pad $\hat{A}_{T+h}$ or ${A}_{T+h}$ with rows and columns of zeros to make their sizes sthe same. 3. Degree cosine similarity measuring the cosine similarity between the degree vectors of $\hat{\mathcal{G}}_{T+h}$ and ${\mathcal{G}}_{T+h}$. Again, we make the vectors compatible by inserting zeros at the end if the lengths differ.
@@ 4 Results
We investigate the performance of the network prediction algorithm on synthetic and real datasets.
@@ 4.1 Synthetic Experiments
For synthetic experiments, we generate a growing sequence of graphs $\left\{\mathcal{G}_{t}\right\}_{t=1}^{T}$ as follows. We start with a 5 node graph $\mathcal{G}_{1}$ and consider a node growth rate of $10\%$, i.e, graph $\mathcal{G}_{t+1}$ has $10\%$ more nodes compared to $\mathcal{G}_{t}$. To generate the graph sequence we consider 2 parameters: proportion of edges removed $\alpha$ and the proportion of edges increased $\beta$. From each $\mathcal{G}_{t}$ we randomly remove $\alpha|E_{t}|$ edges, where $|E_{t}|$ denotes the number of edges in $\mathcal{G}_{t}$. Thus, low values of $\alpha$ indicate a higher proportion of common edges between $\mathcal{G}_{t}$ and $\mathcal{G}_{t+1}$.
The parameter $\beta$ denotes the proportion of edges increased from $\mathcal{G}_{t}$ to $\mathcal{G}_{t+1}$, i.e., $\beta=(|E_{t+1}|-|E_{t}|)/|E_{t}|$. A larger $\beta$ value indicates a higher graph growth in terms of edges. Even though $\alpha$ and $\beta$ measure somewhat similar quantities, the difference is that $1-\alpha$ gives a similarity measure between $\mathcal{G}_{t}$ and $\mathcal{G}_{t+1}$ and $\beta$ gives a growth metric. Thus, they measure different aspects of graph evolution. Both edge deletion and addition are done randomly. Using this setting we predict graphs $\mathcal{G}_{T+h}$ for $T\in\{20,\ldots,24\}$ and $h\in\{1,\ldots,5\}$. We vary the parameters $\alpha$ and $\beta$ such that $\alpha,\beta\in\{0.1,0.2,\ldots,0.9\}$.

### 25 Jul 2024 v2

@@ 3.3 The matrix $S$ and the union graph $\mathcal{G}_{U}$
($\bm{S}$ from a graph sequence $\bm{\{{\mathcal{G}_{t}}\}_{t=1}^{T}}$) For a sequence of graphs $\{{\mathcal{G}_{t}}\}_{t=1}^{T}$ we consider the graph union $\mathcal{G}_{U}=\mathcal{G}_{1}\bigcup\ldots\bigcup\mathcal{G}_{T}$, and add vertices $\hat{n}_{T+h}-n_{T}$, if this quantity is positive. Furthermore, we add edges from new vertices to verticies in $\kappa$ (the $K$ most popular nodes in the union graph $\mathcal{G}_{U}$) and compute matrix $S$.
@@
where $\bm{u}$ denotes an $n(n-1)/2$ vector of edges and $\bm{d}$ denotes an $n$ vector of degree forecasts with $n=\hat{n}_{T+h}$. For all new (unseen) vertices $k$ we use the same upper bound $f_{u}\left(\hat{d}_{k,T+h}\right)$ computed by taking a mean over the node degree for new nodes from the graph time series $\{\mathcal{G}_{t}\}_{t=1}^{T}$ for $t\leq T$. That is, for each graph $\mathcal{G}_{t}$ in the time series, we consider the new nodes at time $t$ and take their average degree. For the remainder of this section we use $n$ to denote $\hat{n}_{T+h}$.
3.4 Coefficients $\xi_{ij}$
Of the many graph forecasts possible, we want to select a solution that depicts the most likely graph at time $T+h$. However, the term “most likely” can be subjective. Our definition of most likely involves maximizing the function $\sum_{ij}\xi_{ij}\hat{e}_{ij,T+h}$ where $\xi_{ij}$ denote the coefficients of the optimization function. The coefficients satisfy $0\leq\xi_{ij}\leq 1$ with higher values given to edges more likely to be present in $\mathcal{G}_{T+h}$. We state the complete optimization problem below:
$\max\sum_{i,j\in V_{T+h}}\xi_{ij}\hat{e}_{ij,T+h}\,,\text{with}~{}~{}\bm{0}\leq S\bm{u}\leq f_{u}\left(\bm{d}\right)\,,$ (3)
@@ 3.4 Coefficients $\xi_{ij}$
    This coefficient function is naive as it considers any edge to be equally likely including edges that have never been present in the past. This function may be useful for instances with limited history or when we want to completely disregard the history of a network when forecasting edges at time $T+h$.
@@
    Proportional coefficients (C3): Here we use
    $\xi_{ij}=\begin{cases}\xi_{ij}^{\text{exist}}&\text{for existing vertices }i\text{ and }j\\ \xi_{ij}^{\text{new}}&\text{ for new vertices }j\end{cases}\,.$ (5)
    For existing vertices $i$ and $j$ we have
    $\xi_{ij}^{\text{exist}}=\frac{\sum_{t\leq T}\bm{1}_{e_{ij,t}\in\mathcal{G}_{t}}}{\max_{i,j}\sum_{t\leq T}\bm{1}_{e_{ij,t}\in\mathcal{G}_{t}}}$ (6)
    where $\bm{1}$ is the indicator function
    $\bm{1}_{e_{ij,t}\in\mathcal{G}_{t}}=\begin{cases}1\quad\text{if }e_{ij,t}\in\mathcal{G}_{t}\\ 0\quad\text{otherwise}\end{cases}\,.$ (7)
    The term $\sum_{t\leq T}\bm{1}_{e_{ij,t}\in\mathcal{G}_{t}}$ counts the number of times edge $e_{ij}$ is present in the graph time series. Thus, for existing vertices at time $T$, we take the proportion of time points when edge $e_{ij}$ is present. For new vertices $j$ connecting to existing vertices $i$ we have
    $\xi_{ij}^{\text{new}}=Q\left(\frac{d_{i,U}}{\sum_{k\in\kappa}d_{k,U}}\right)\,,$
    where $Q$ is the quantile function with respect to the distribution of $\xi_{ij}^{\text{exist}}$ where both $i$ and $j$ are existing vertices. The term $d_{i,U}$ denotes the degree of node $i$ in the union graph. The $\frac{d_{i,U}}{\sum_{k\in\kappa}d_{k,U}}$th quantile used. For new vertices, we assign a larger weight if it connects to a more popular existing node. Recall that edges are added to $\mathcal{G}_{U}$ connecting new nodes to nodes in $\kappa$, the $K$ most popular nodes in $\mathcal{G}_{U}$. The weights for these edges depend on the degree of the node $v_{i}\in\kappa$. We want the distribution $\xi_{ij}^{\text{exist}}$ to match with $\xi_{ij}^{\text{new}}$. To achieve this we take quantiles proportional to $d_{i,U}$ from $\xi_{ij}^{\text{exist}}$.
@@
    Linear-decay coefficients (C4):
    $\xi_{ij}=\begin{cases}\frac{\sum_{t\leq T}t\cdot\bm{1}_{e_{ij,t}\in\mathcal{G}_{t}}}{\max_{i,j}\sum_{t\leq T}t\cdot\bm{1}_{e_{ij,t}\in\mathcal{G}_{t}}}&\text{for existing vertices }i\text{ and }j\\ Q\left(\frac{d_{i,U}}{\sum_{k\in\kappa}d_{k,U}}\right)&\text{if }j\text{ is a new vertex}\end{cases}$ (8)
    This is a linearly weighted version of C3 coefficients with higher weights given to more recent time points. Thus, the coefficients linearly decay with time.
  5. 5.
    Harmonic-decay coefficients(C5):
    $\xi_{ij}=\begin{cases}\frac{\sum_{t\leq T}\left(\frac{1}{T-t+1}\cdot\bm{1}_{e_{ij,t}\in\mathcal{G}_{t}}\right)}{\max_{i,j}\sum_{t\leq T}\left(\frac{1}{T-t+1}\cdot\bm{1}_{e_{ij,t}\in\mathcal{G}_{t}}\right)}&\text{for existing vertices }i\text{ and }j\\ Q\left(\frac{d_{i,U}}{\sum_{k\in\kappa}d_{k,U}}\right)&\text{if }j\text{ is a new vertex}\end{cases}$ (9)
    This too is a weighted version of C3 with weights for edges of graph $\mathcal{G}_{t}$ multiplied by $\frac{1}{T-t+1}$. Thus, edges in $\mathcal{G}_{T}$ would get initially multiplied by 1, and edges in $\mathcal{G}_{T-1}$ would get multiplied by 1/2 and so on. The coefficients are harmonically decaying in time.
  6. 6.
    Last seen graph coefficients (C6):
    $\xi_{ij}=\begin{cases}\bm{1}_{e_{ij,T}\in\mathcal{G}_{t}}&\text{for existing vertices }i\text{ and }j\\ Q\left(\frac{d_{i,U}}{\sum_{k\in\kappa}d_{k,U}}\right)&\text{if }j\text{ is a new vertex}\end{cases}$ (10)
    In this scheme we assign 1 to edges existing in the last seen graph $\mathcal{G}_{T}$ and 0 to other edges in the union graph $\mathcal{G}_{U}$. For edges from new vertices we use the same quantile function as before.
The above coefficient schemes can be seen as specific instances of a generalized version given by
$\xi_{ij}=\begin{cases}f\left(e_{ij}\right)&\text{ for existing vertices }i,j\\ \mu_{i}&\text{ if }v_{i}\text{ is existing and }v_{j}\text{ is new }\\ \end{cases}$ (11)
where $f\left(e_{ij}\right)$ is some prediction of $e_{ij}$ and $\mu_{i}$ depends on existing node $v_{i}$.
Considering the coefficients $\{\xi_{ij}\}_{v_{i},v_{j}\in V_{t+H}}$, the question arises why these alone cannot be used to describe the graph $\mathcal{G}_{T+h}$. We emphasize that $\xi_{ij}$ is computed by considering individual edges $e_{ij}$ or nodes $v_{i},v_{j}$, without taking the network structure into account. Thus, $\xi_{ij}$ is not a good approximation for $e_{ij,T+h}$ and we do not recommend it to be used by itself to construct $\mathcal{G}_{T+h}$. However, $\xi_{ij}$ provides good value as coefficients of the optimization problem described in equation (3).
3.5 The optimization problem
We specify two versions of the optimization problem. The first version has the individual degree constraints discussed so far, and the second version has an additional constraint on total edges.
3.5.1 Formulation 1 (F1)
Consider the following optimization problem:
$\max\sum_{i,j\in V_{T+h}}\xi_{ij}\hat{e}_{ij,T+h}\,,\text{with}~{}~{}\bm{0}\leq S\bm{u}\leq f_{u}\left(\bm{d}\right)\,,$ (12)
with $\bm{u}=\{u_{\ell}\}_{\ell=1}^{n(n+1)/2}$ and $u_{\ell}\in\{0,1\}$ where each inequality in (12) describes a constraint for the degree upper bound of a single vertex $v_{i}$ given by
$\sum_{j}\hat{e}_{ij,T+h}\leq f_{u}\left(\hat{d}_{i,T+h}\right)\,.$
So far each constraint considers the degree of a single node. However, as we are considering upper bounds we may add a lot more edges to the graph. By constraining the total edges using its prediction can help us to get a more realistic graph. We note that this is a global constraint affecting all nodes. Next we use this additional, global constraint in our optimization.
3.5.2 Formulation 2 (F2)
Consider the following optimization problem:
@@ 3.5.2 Formulation 2 (F2)
$\text{with}\qquad\bm{0}$ $\leq S\bm{u}\leq f_{u}\left(\bm{d}\right)\,,$
@@
The proof is given in the Appendix.
3.6 $\hat{\mathcal{G}}_{T+h}$ and the forecast distribution of graphs
By solving the optimization problem described in Formulations 1 or 2 we obtain a possible graph at time $T+h$, which we denote by $\hat{\mathcal{G}}_{T+h}$. This graph is a forecast and thus is an element of a forecast distribution. In fact it was obtained by making certain choices on $n$, (used to denote $\hat{n}_{T+h}$) and $f_{u}(\bm{d})$.
Figure 2 illustrates the forecast graph distribution in which $\hat{\mathcal{G}}_{T+h}$ is a function of forecast nodes $\hat{n}_{T+h}$ and $f_{u}(\bm{d})$. If we explore the forecasts for different values of $\hat{n}_{T+h}$ and different upper bounds $u$, we get different graphs. Let $\gamma$ denote the quantile used to obtain $\hat{n}_{T+h}$ and $u$ denote the quantiles in $f_{u}(\bm{d})$. Then we can think of the graph $\hat{\mathcal{G}}_{T+h}$ as belonging to a graph distribution $\mathfrak{G}$ having parameters $\gamma$ and $u$:
@@ 3.6 $\hat{\mathcal{G}}_{T+h}$ and the forecast distribution of graphs
While these are explicit parameters, the coefficient scheme used to obtain $\xi_{ij}$, the methods used for forecasting individual time series and their hyperparameters can also be considered as part of a wider parameter pool.
Figure 2: A simple illustration of the forecast graph distribution in terms of $\gamma$ and $u$. The parameter $\gamma$ gives different $\hat{n}_{T+h}$ values. The 2 graphs on the left have 4 nodes each, corresponding to a single $\gamma$ (or $\hat{n}_{T+h}$) value, but has a different number of edges corresponding to different $u$ values. Similarly, the two graphs on the right have 7 nodes corresponding to a higher value of $\gamma$ with the rightmost graph having more edges due to a higher $u$.
@@ 4 Results
We investigate the performance of the proposed network prediction algorithm on synthetic and real datasets. For all experiments we consider a time series of graphs $\{\mathcal{G}_{t}\}_{t=1}^{T}$ and forecast graphs at future time steps for $h\in\{1,\ldots,5\}$ using Formulation 2 (F2) with harmonic (C5) and last seen (C6) coefficient schemes. The R package netseer has functionality for both F1 and F2 formulations and other coefficient schemes. For comparison we use the last seen graph $\mathcal{G}_{T}$.
4.0.1 Evaluation Metrics:
We use the following evaluation metrics: (1) Node error $\frac{|\hat{n}_{T+h}-n_{T+h}|}{n_{T+h}}$ gives the absolute error ratio of the number of vertices in the predicted graph compared to the actual graph; (2) Edge error $\frac{|\hat{m}_{T+h}-m_{T+h}|}{m_{T+h}}$ gives the absolute error ratio of the number of edges in the predicted graph compared to the actual graph; (3) Edge density error ${|\hat{\text{dens}}(T+h)-\text{dens}(T+h)|}$ gives the absolute error of edge densities between the predicted graph and the actual graph. As the edge density is between 0 and 1 we take the absolute difference, not a ratio. For all 3 evaluation metrics, a lower absolute error is preferred.
@@ 4.1 Synthetic Experiments
For synthetic experiments, we generate a growing sequence of graphs $\left\{\mathcal{G}_{t}\right\}_{t=1}^{T}$ using Preferential Attachment (PA) (Barabási & Albert, 1999). The PA model considers nodes connecting to more connected nodes with a higher probability. The linear PA model specifies the probability $\Pi(k)$ of a new node connecting to node $i$ with degree $k_{i}$ to be
$\Pi(k_{i})=\frac{k_{i}}{\sum_{i}k_{i}}\,.$
At each time step, the PA model adds a new node with $s$ edges. After $t$ timesteps the network has $n_{t}=t+s_{0}$ nodes and $m_{t}=s_{0}+ts$ edges where at time $t=0$ the network has $s_{0}$ nodes and edges.
As the first experiment we consider a sequence of 30 PA networks with $s=10$, $n_{1}=50$ and $n_{30}=350$. At each timestep, 5 nodes are added and each node connects to 10 other nodes. Using the first 25 networks, we forecast the next 5 networks, i.e., we set $T=25$ and forecast $\hat{\mathcal{G}}_{T+h}$ for $h\in\{1,\ldots,5\}$. In this experiment $E_{t}\subset E_{t+1}$ as edges are added to the existing network. We repeat the experiment 10 times with different seeds to account for randomness.
The second experiment includes edge deletion. After adding new nodes and new edges as before, at each time step we randomly delete $10$ edges from the network. As such $E_{t}\not\subset E_{t+1}$. Again, we repeat the computation 10 times to account for randomness.
Table 1: Experiment 1 and 2 Results. Mean and standard deviation in brackets of 10 iterations.
Exp. Metric Coef. $h=1$ $h=2$ $h=3$ $h=4$ $h=5$
1 Node C5 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
Err. C6 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
LS 0.0286 (0) 0.0556 (0) 0.0811 (0) 0.1053 (0) 0.1282 (0)
Edge C5 0 (0) 0.0077 (0.0014) 0.0196 (0.0016) 0.0318 (0.0024) 0.0435 (0.0025)
Err. C6 0 (0) 0.004 (7e-04) 0.0099 (9e-04) 0.016 (0.0012) 0.0218 (0.0013)
LS 0.0295 (0) 0.0573 (0) 0.0836 (0) 0.1084 (0) 0.1319 (0)
Dens C5 0 (0) 0.0077 (0.0014) 0.0196 (0.0016) 0.0318 (0.0024) 0.0435 (0.0025)
Err. C6 0 (0) 0.004 (7e-04) 0.0099 (9e-04) 0.016 (0.0012) 0.0218 (0.0013)
LS 0.0286 (0) 0.0572 (0) 0.0858 (0) 0.1144 (0) 0.143 (0)
2 Node C5 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
Err. C6 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
LS 0.0286 (0) 0.0556 (0) 0.0811 (0) 0.1053 (0) 0.1282 (0)
Edge C5 6e-04 (0) 6e-04 (2e-04) 0.0087 (0.0028) 0.0205 (0.0028) 0.0321 (0.0028)
Err C6 0 (0) 6e-04 (7e-04) 0.0062 (0.0013) 0.0119 (0.0013) 0.0178 (0.0013)
LS 0.0297 (0) 0.0576 (0) 0.084 (0) 0.109 (0) 0.1326 (0)
Dens. C5 6e-04 (0) 6e-04 (2e-04) 0.0087 (0.0028) 0.0205 (0.0028) 0.0321 (0.0028)
Err. C6 0 (0) 6e-04 (7e-04) 0.0062 (0.0013) 0.0119 (0.0013) 0.0178 (0.0013)
LS 0.0284 (0) 0.0568 (0) 0.0853 (0) 0.1137 (0) 0.1421 (0)